Prove or disprove that the following vector function is discontinuos
r(t)=(cost)i+(sint)j+[t] k
1
Expert's answer
2020-12-10T14:21:55-0500

This is from "Multivariable Calculus, Concepts and Contexts" by Stewart.

He says "The parametric equations for this curve are: "x=cos(t), y=sin(t), z=t"


We know from trigonometry that "cos^2t+sin^2t=1" , and you know from the parametric equation of the curve that ."x=cost,y=sint". So for any point of the curve, it is true that "x^2+y^2=1"


Just notice that you dropped the equality "z=t", meaning that you get more "(x,y,z)"

values that satisfy the equalities: you obtain a superset of the curve points. Indeed, the parametric equation corresponds to a linear entity (1D), while "x^2+y^2=1" defines a surface (2D), which contains the curve.


Therefore the vector function is continuos